OpenStudy (ttop0816):

is my answer correct?? Find the arc length intercepted by a central angle of pi/2 radians in a circle whose radius is 24 inches. 48π 24π 12π <--- i chose this!

OpenStudy (phi):

arc length is angle (in radians) * radius in this case \[ \frac{\pi}{2} \cdot 24 = 12 \pi \text{ inches} \]

OpenStudy (ttop0816):

@phi thank you! what about this question?? ): In what quadrant does the angle -8pi/5 terminate? I <--- i chose this! II III IV